# A Learning-Augmented Overlay Network

### v2

@@ 1.2 Contributions
As our third contribution, we present a learning-augmented overlay algorithm, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform algorithm with a predicted SLN to strike consistency and robustness guarantees (7). For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $[0,1]$ uniformly at random and adds it to the integer heights to obtain its final heights as a C-SLN; it may then join the network using the join() procedure. We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $O(\log^{2}n)$ factor away from the optimum static SLN in case of adversarial predictions ($O(\log^{2}n)$-robust).
@@
Figure 1: A SLN with eight nodes. The node ids range from $1$ to $8$. The double-arrowed, horizontal lines are the edges of the SLN.
@@ 3 Optimum Static Skip List Network
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$. Let $nl$, $1\leq nl\leq l$, denote the leftmost node id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost node id, such that $\forall v\in[r,nr],\,h_{v}\leq h$. Finally, let $\textsc{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$.
@@
Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$). The minimum cost for the solution for any $nl$ and $nr$ is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
@@ 4.2 Uniform C-SLN
Let $u$ and $v$ be two nodes. We note $X$ the random variable equal to the routing path length between $u$ and $v$. We note $X_{\text{up}}$ the number of height increases along the path, likewise $X_{\text{down}}$ is the number of path decreases. It holds that $X=X_{\text{up}}+X_{\text{down}}$. We will give an upper bound of $X_{\text{up}}$, the same upper bound will hold for $X_{\text{down}}$ by symmetry. Let $w$ be a node between $u$ and $v$ in the SLN, we note $1_{w}$ the indicator random variable equal to $1$ if (1) $w$ is part of the path between $u$ and $v$, and (2) if the height increases when hoping to $w$. It holds $X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$.
@@
Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ . Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 3).
@@
For Uniform, the expected degree of any node is lower than $4$.
@@
Let $u$, $w$ and $w^{\prime}$ be three different nodes. We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively. We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 4.
@@ 5 Learning-Augmented C-SLN
As the considered algorithms has a logarithmic maximum degree with high probability, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n). Let $u$ be a node. We first cover the case when the maximum degree is lower than $\ln^{c}n$, all the considered probabilities and expected values are therefore conditional. In this case, we consider the disk centered around $u$ with smallest radius such that it contains no less than half of the nodes. The maximum distance between $u$ and a node in that disk is no lower than $O(\log n/\log\log n)$ as the network has a polylogarithmic maximum degree — this result can be easily obtained by upper bounding all the degrees by $O(\log^{c}n)$. There exists another node $v$ such that $v$ is at the border or outside of the disc with probability at least $1/2$. We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$. An optimal SLN algorithm can clearly pay a cost of $1$, the considered SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
@@ 6 Experimental Evaluation
  • Adversarial Noise: This model simulates the effect of the presence of malicious nodes in the network. For every network node, the height value assigned to it by the Skip List algorithm is reset to $1$, with probability $\eta$.
@@
In the Adversarial Noise model, adversarial agents reset their own node heights to a baseline value of $1$, with probability $\eta$, eliminating structural (SLN) shortcuts and increasing the total routing work (communication cost). The objective of this experiment is to demonstrate the efficacy of a simple countermeasure: by increasing each node’s height value with a small randomized value, we can effectively safeguard the network’s performance against such adversarial intent.

### v3

@@ 1.2 Contributions
As our third contribution, we present a learning-augmented overlay algorithm, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform algorithm with a predicted SLN to strike consistency and robustness guarantees (7). For each node, LASLiN takes as prediction an integer between 11 and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $[0,1]$ uniformly at random and adds it to the integer heights to obtain its final heights as a C-SLN; it may then join the network using the join() procedure. We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $O(\log^{2}n)$ factor away from the optimum static SLN in case of adversarial predictions ($O(\log^{2}n)$-robust).
@@
Figure 1: A SLN with eight nodes. The node ids range from 11 to 88. The double-arrowed, horizontal lines are the edges of the SLN.
@@ 3 Optimum Static Skip List Network
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$. Let nlnl, $1\leq nl\leq l$, denote the leftmost node id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let nrnr, $r\leq nr\leq n$, denote the rightmost node id, such that $\forall v\in[r,nr],\,h_{v}\leq h$. Finally, let $\textsc{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$.
@@
Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$). The minimum cost for the solution for any nlnl and nrnr is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
@@ 4.2 Uniform C-SLN
Let $u$ and $v$ be two nodes. We note $X$ the random variable equal to the routing path length between $u$ and $v$. We note $X_{\text{up}}$ the number of height increases along the path, likewise $X_{\text{down}}$ is the number of path decreases. It holds that $X=X_{\text{up}}+X_{\text{down}}$. We will give an upper bound of $X_{\text{up}}$, the same upper bound will hold for $X_{\text{down}}$ by symmetry. Let $w$ be a node between $u$ and $v$ in the SLN, we note $1_{w}$ the indicator random variable equal to 11 if (1) $w$ is part of the path between $u$ and $v$, and (2) if the height increases when hoping to $w$. It holds $X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$.
@@
Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ . Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to 11 if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 3).
@@
For Uniform, the expected degree of any node is lower than 44.
@@
Let $u$, $w$ and $w^{\prime}$ be three different nodes. We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to 11 if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively. We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 4.
@@ 5 Learning-Augmented C-SLN
As the considered algorithms has a logarithmic maximum degree with high probability, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n). Let $u$ be a node. We first cover the case when the maximum degree is lower than $\ln^{c}n$, all the considered probabilities and expected values are therefore conditional. In this case, we consider the disk centered around $u$ with smallest radius such that it contains no less than half of the nodes. The maximum distance between $u$ and a node in that disk is no lower than $O(\log n/\log\log n)$ as the network has a polylogarithmic maximum degree — this result can be easily obtained by upper bounding all the degrees by $O(\log^{c}n)$. There exists another node $v$ such that $v$ is at the border or outside of the disc with probability at least $1/2$. We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$. An optimal SLN algorithm can clearly pay a cost of 11, the considered SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
@@ 6 Experimental Evaluation
  • Adversarial Noise: This model simulates the effect of the presence of malicious nodes in the network. For every network node, the height value assigned to it by the Skip List algorithm is reset to 11, with probability $\eta$.
@@
In the Adversarial Noise model, adversarial agents reset their own node heights to a baseline value of 11, with probability $\eta$, eliminating structural (SLN) shortcuts and increasing the total routing work (communication cost). The objective of this experiment is to demonstrate the efficacy of a simple countermeasure: by increasing each node’s height value with a small randomized value, we can effectively safeguard the network’s performance against such adversarial intent.